40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1123     Entry time: Wed Mar 20 20:44:32 2013
Author: Evan 
Type: DailyProgress 
Category: optic 
Subject: Added Faraday isolator 

Today I installed the Faraday isolator after the PMC. Tara and I then spent some time trying to figure out why the PDH error signal suddenly had a huge DC offset (it was because I accidentally knocked the angle control on one of the HWP mounts while installing the FI beam dump). Before installing the FI, we had observed that the loop oscillates noticeably at about 100 kHz and had hoped it was caused by back-reflection into the laser (which the FI would fix). Installing the FI seems to have no effect on the oscillations. After installing the FI I adjusted the HWP immediately following and retuned the phasing of the PDH loop by adding some extra cable to the PD SMA input. I've attached a picture showing the sweeps of the cavity refl response and PDH error signal, and a picture showing the oscillations when the loop is engaged.

I tried minimizing the rejected light out of the FI to optimize the angle of the QWP directly in front of the cavity, but this light appears to be dominated by reflections other than those off of the cavity. The rejected light consists of two distinct spots which can be seen with an IR card. I think one of them is a reflection from the lens immediately following the FI, and the other is a reflection from the 14.75 MHz EOM.

I also tried improving the mode matching into the cavity by adjusting the two lenses on translation stages, but I didn't get much improvement. Right now we are at just under 50% coupling.
Attachment 1: cavity_scan.jpg  109 kB  | Hide | Hide all
cavity_scan.jpg
Attachment 2: loop_oscillation.jpg  109 kB  | Hide | Hide all
loop_oscillation.jpg
ELOG V3.1.3-